# package/before/amsmath

### 14 Dec 2023 v1

@@ 2.3 Mean Variance Criteria of Performance Aware Agents
We can extend our framework to consider agents that are aware about their performance relative to others. We refer to these agents as performance conscious agents of degree $\phi$. More formally, for agent $k$, let the market average wealth be given by Guan and Hu [2022]:
@@
Then for $\phi\in[0,1)$ we let $\hat{X}(t)=X_{k}(t)-\phi\bar{X}(t)$ and the mean-variance criteria becomes Guan and Hu [2022]:
@@
where $\gamma_{k}>0$ is the risk aversion parameter of each agent $k$. Note that by setting $\phi=0$ in (10), the cost function reduces to (4). The solution to the $K$-agent problem results in an optimal control Guan and Hu [2022]. The Nash equilibrium has been shown in Guan and Hu [2022] and is given by the following definition:
@@
Optimal control of time consistent mean-variance problem for $K$-agents Guan and Hu [2022]. Let $v_{k}\in\mathcal{V}_{k}$ be the admissible control $\forall k=1,...,K$. A vector $(v_{1}^{*},v_{2}^{*},...,v_{K}^{*})$ is said to be a time-consistent optimal strategy, any fixed initial state $(t,x,y)\in[0,T]\times\mathbb{R}\times\mathbb{R}$ and a fixed number $h>0$, a new vector $(v_{1}^{h},v_{2}^{h},...,v_{K}^{h})$ defined by:
@@
The time-consistent solution satisfies the conditions of a subgame perfect Nash equilibrium Van Staden et al. [2018]. We remark that in this paper, we use the term optimal control instead of equilibrium control in order to be consistent with other numerical approaches in the literature Van Staden et al. [2018]; Wang and Forsyth [2011]; Bjork et al. [2017] .
@@ 3 Backward Stochastic Differential Equation (BSDE) Representation of the HJB Equation
In this section, we present the optimal trade execution problem using an auxiliary HJB formulation and its BSDE. The HJB formulation has been shown to have a viscosity solution Aivaliotis and Veretennikov [2018]. We also show that when there is no price impact, the formulation reduces to the optimal investment problem. We reformulate the HJB to the BSDE in order to solve the optimal trade execution problem as a stochastic optimization problem which is more suitable for applications of neural networks.
@@ 3.1 HJB Formulation
The value function $u$ can be solved by reformulating the problem as an auxiliary problem Wang and Forsyth [2011]. However, the auxiliary problems do not permit a viscosity solution. To overcome this, we can rewrite the variance term such that Aivaliotis and Veretennikov [2018]:
@@
where $U(S,b,\alpha,t)$ is the value function of the Markovian control problem given by Aivaliotis and Veretennikov [2018]:
@@
We extend this formulation to a performance conscious agent $k$ Guan and Hu [2022] by replacing $X(T)$ with $\hat{X}(T)=X(T)-\phi\bar{X}(T)$. We formally extend (13) to $K$ agents and $d$ assets in the following Lemma:
@@
$+\sum_{i=1}^{d}v_{i,k}\frac{\partial U_{k}}{\partial\alpha_{i,k}}-v_{k}\sum_{i=1}^{d}(S_{i}f(v_{i,k}))\frac{\partial U_{k}}{\partial b_{k}}\}=0.$ (14)
@@
Let $X(t)\equiv(S,b,A,t)$, where $S=\sum_{i=1}^{d}S_{i}(t)$ and $A_{k}=\sum_{i=1}^{d}\alpha_{i,k}(t)$. Let $\Delta t$ be the timestep size, $\Delta S=S(t+\Delta t)-S(t),\Delta b=b(t+\Delta t)-b(t)$, and $\Delta A=A(t+\Delta t)-A(t)$, then by the optimality of the control and law of iterated expectation for each $k$ the value function of the auxiliary problem is given by:
$U_{k}=U_{k}(S,b_{k},A_{k},t)=\sup_{v\in\mathcal{V}}\mathbb{E}[U_{k}(S+\Delta S,b_{k}+\Delta b_{k},A_{k}+\Delta A_{k},t+\Delta t)].$ (15)
Let $\Delta U_{k}=U_{k}(S+\Delta S,b_{k}+\Delta b_{k},A_{k}+\Delta A_{k},t+\Delta t)-U_{k}(S,\hat{b},\hat{\alpha},t)$, then (15) can be written as
@@
Applying Ito’s Lemma on $\Delta U$, we get:
$\mathbb{E}[\Delta U]=\mathbb{E}\left[\frac{\partial U}{\partial t}\Delta t+\frac{\partial U}{\partial S}\Delta S+\frac{1}{2}\frac{\partial^{2}U}{\partial S^{2}}(\Delta S)^{2}+\frac{\partial U}{\partial b}\Delta b+\frac{\partial U}{\partial A}\Delta A\right],$
@@
$\mathbb{E}[\Delta S]$ $=\sum_{i=1}^{d}(\mu_{i}-r+g(\sum_{k=1}^{K}v_{i,k}))S_{i}\Delta t,$
$\mathbb{E}[\Delta\hat{b}]$ $=(r\hat{b}-v_{k}\sum_{i=1}^{d}(S_{i}f(v_{i,k})))\Delta t,$
$\mathbb{E}[\Delta\hat{A}]$ $=\sum_{i=1}^{d}v_{i,k}\Delta t.$
The expectations of $\Delta S,\Delta b$ and $\Delta A$ are straight forward. The expectation of $(\Delta S)^{2}$ requires more thought. Applying Ito’s lemma on $(\Delta S)^{2}$, the expected value is given by:
$\mathbb{E}[(\Delta S)^{2}]=\sum_{i,j=1}^{d}\rho_{i,j}\sigma_{i}\sigma_{j}(S_{i})^{2}\Delta t.$
Expanding (16) and evaluating $\mathbb{E}[\Delta S],\mathbb{E}[(\Delta S)^{2}],\mathbb{E}[\Delta b],\mathbb{E}[\Delta A]$ gives us:
$\left[\frac{\partial U_{k}}{\partial t}+\mathcal{L}_{k}U_{k}+rb_{k}\frac{\partial U_{k}}{\partial b_{k}}\right.$ $+\sup_{v\in\mathcal{V}}\left\{\sum_{i=1}^{d}g(\sum_{k=1}^{K}v_{i,k})S_{i}\frac{\partial U_{k}}{\partial S}\right.$
$\left.\left.+\sum_{i=1}^{d}v_{i,k}\frac{\partial U_{k}}{\partial\alpha_{i,k}}-v_{k}\sum_{i=1}^{d}(S_{i}f(v_{i,k}))\frac{\partial U_{k}}{\partial b_{k}}\right\}\right]\Delta t=0,$
where
$\mathcal{L}_{k}U=\sum_{i=1}^{d}(\mu_{i}-r)S_{i}\frac{\partial U}{\partial S_{i}}+\sum_{i,j=1}^{d}\rho_{i,j}\sigma_{i}\sigma_{j}(S_{i})^{2}\frac{\partial^{2}U}{\partial S_{i}^{2}}.$
Diving out the $\Delta t$ and bringing the supremum inside the expectation gives us the extended HJB equation. ∎
@@
For the special case of $\kappa_{\tau}=0,\kappa_{p}=0$, i.e. there is no price impact, the optimal trade execution reduces to the optimal asset allocation problem. The analytical solution to the optimal asset allocation problem for time-consistent mean-variance problem is known and has been solved in the linear-quadratic (LQ) sense Zhou and Li [2000]; Li and Ng [2000].
@@ 3.2 BSDE Formulation
A stochastic approach has an advantage in that the HJB can be extended to include multiple assets and multiple agents. This is computationally infeasible with the PDE approach even with the dimensional reduction techniques Forsyth [2011]. This is because each PDE for variance and wealth requires a three dimensional grid even with one asset. Though dimensional reduction can be used, this is does not apply for the general performance conscious agent due to the averaging term. We drop the explicit dependence on $t$ to keep the notation concise. To solve (14), we formulate the BSDE for $U_{k}$ using the following lemma.
@@
Given the control $v_{i,k}$, the BSDE that satisfy viscosity solution for (14) is given by:
@@
$\frac{\partial U_{k}}{\partial t}+\mathcal{L}_{k}U_{k}$ $=-rb_{k}\frac{\partial U_{k}}{\partial b_{k}}+v_{k}\sum\limits_{i=1}^{d}f(v_{i,k})\frac{\partial U_{k}}{\partial b_{k}}$
$-\sum\limits_{i=1}^{d}(v_{i,k})\frac{\partial U_{k}}{\partial\alpha_{i,k}}-\sum\limits_{i=1}^{d}(g(\sum_{k=1}^{K}v_{i,k}))S_{i}\frac{\partial U_{k}}{\partial S_{i}}.$
@@
The proof is from the direct application of the BSDE formulation given in Pham [2014]. Let $x_{i,k}=\{S_{i},b_{k},\alpha_{i,k}\}$, for each agent $k$. Given the control $v_{i,k}$ and a semi-linear PDE given by:
$\frac{\partial U_{k}}{\partial t}+\mathcal{L}_{k}U_{k}+\sum_{i=1}^{d}f(x_{i,k},v_{i,k},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)=0,$
@@
$dU_{k}=-\sum_{i=1}^{d}f(x_{i,k},v_{i,k},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)dt+\sum_{i=1}^{d}\sigma_{i}\frac{\partial U_{i,k}}{\partial x_{i,k}}dW_{i},$
from the semi-linear PDE we can express $-f(x_{i,k},v_{i,k},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)$ as:
$-\sum_{i=1}^{d}f(x_{i,k},v_{i,k},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)=\frac{\partial U_{k}}{\partial t}+\mathcal{L}_{k}U_{k}.$
@@
$-\sum_{i=1}^{d}f(x_{i,k},v_{i,k},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)$ $=-rb_{k}\frac{\partial U_{k}}{\partial b_{k}}+v_{k}\sum\limits_{i=1}^{d}f(v_{i,k})\frac{\partial U_{k}}{\partial b_{k}}$
$-\sum\limits_{i=1}^{d}(v_{i,k})\frac{\partial U_{k}}{\partial\alpha_{i,k}}-\sum\limits_{i=1}^{d}(g(\sum_{k=1}^{K}v_{i,k}))S_{i}\frac{\partial U_{k}}{\partial S_{i}}.$
The significance of the system of BSDEs is two fold. One, the system of BSDEs does not require us to compute the Hessian of the extended HJB equation. This a big advantage as it allows us to avoid computing the second derivative on $U$ since computation of the second derivative can be very costly for high dimensions. This reduces the computational complexity of the problem. The second significance of the system of BSDEs is that it allows us to solve (17) in a least squares sense which is a recent technique used in high-dimensional option pricing Han and E [2016]; E et al. [2017]; Hure et al. [2020]; Chen and Wan [2021]; Na and Wan [2023]. This allows us to avoid using numerical solutions to the PDE and overcome the issue with dimensionality. Note, we can replace $X$ in (11) and (12) with $\hat{X}$ for the case where the agent is performance conscious, i.e. $\phi>0$.
@@ 3.3 Discretization of BSDE
$(S_{i})_{m}^{n+1}$ $=(1+(\mu_{i}-r+g(\sum_{k=1}^{K}(v_{i,k})_{m}^{n}))\Delta t)(S_{i})_{m}^{n}+\sigma_{i}(S_{i})_{m}^{n}(\Delta W_{i})_{m}^{n},$ (18)
@@
The number of shares in stocks, $\alpha(t)$, is updated using (9). Thus the discretized dynamics of $\alpha(t)$ is given by:
@@
Let $\mathcal{V}:=[v_{\min},v_{\max}]$ be the space of bounded trade rates. Then $\hat{\mathcal{V}}$ is the discretization of $\mathcal{V}$ into $C$ number of discrete points. Equations (18), (19), and (20) are simulated forward in time over a subset of discretized control values $\hat{\mathcal{V}}\subset\mathcal{V}$. The discrete set of control points does not need to be too dense and we restrict it to the discrete set shown in Forsyth [2011].

### 19 Mar 2024 v2

@@ 2.3 Mean Variance Criteria of Performance Aware Agents
We can extend our framework to consider agents that are aware about their performance relative to others. We refer to these agents as performance conscious agents of degree $\phi$. More formally, for agent $k$, let the market average wealth be given by [Guan and Hu 2022]:
@@
Then for $\phi\in[0,1)$ we let $\hat{X}(t)=X_{k}(t)-\phi\bar{X}(t)$ and the mean-variance criteria becomes [Guan and Hu 2022]:
@@
where $\gamma_{k}>0$ is the risk aversion parameter of each agent $k$. Note that by setting $\phi=0$ in (8), the cost function reduces to (4). The solution to the $K$-agent problem results in an optimal control [Guan and Hu 2022]. The Nash equilibrium has been shown in [Guan and Hu 2022] and is given by the following definition:
@@
Optimal control of time consistent mean-variance problem for $K$-agents [Guan and Hu 2022]. Let $v_{k}\in\mathcal{V}_{k}$ be the admissible control $\forall k=1,...,K$. A vector $(v_{1}^{*},v_{2}^{*},...,v_{K}^{*})$ is said to be a time consistent optimal strategy, any fixed initial state $(t,x,y)\in[0,T]\times\mathbb{R}\times\mathbb{R}$ and a fixed number $h>0$, a new vector $(v_{1}^{h},v_{2}^{h},...,v_{K}^{h})$ defined by:
@@
The time consistent solution satisfies the conditions of a subgame perfect Nash equilibrium [Van Staden et al. 2018]. We remark that in this paper, we use the term optimal control instead of equilibrium control in order to be consistent with other numerical approaches in the literature [Van Staden et al. 2018; Wang and Forsyth 2011; Bjork et al. 2017] .
@@ 3 Backward Stochastic Differential Equation (BSDE) Representation of the HJB Equation
In this section, we present the optimal trade execution problem using an auxiliary HJB formulation and its BSDE. The HJB formulation has been shown to have a viscosity solution [Aivaliotis and Veretennikov 2018]. We also show that when there is no price impact, the formulation reduces to the optimal investment problem. We reformulate the HJB to the BSDE in order to solve the optimal trade execution problem as a stochastic optimization problem which is more suitable for applications of neural networks.
@@ 3.1 HJB Formulation
The value function $u$ can be solved by reformulating the problem as an auxiliary problem [Wang and Forsyth 2011]. However, the auxiliary problems do not permit a viscosity solution. To overcome this, we can rewrite the variance term such that [Aivaliotis and Veretennikov 2018]:
@@
where $U(S,b,\alpha,t)$ is the value function of the Markovian control problem given by [Aivaliotis and Veretennikov 2018]:
@@
We extend this formulation to a performance conscious agent $k$ [Guan and Hu 2022] by replacing $X(T)$ with $\hat{X}(T)=X(T)-\phi\bar{X}(T)$. We formally extend (11) to $K$ agents and $d$ assets in the following Lemma:
@@
Let the generator $\mathcal{L}U_{k}$ be defined as:
$\mathcal{L}U_{k}=\sum_{i=1}^{d}(\mu_{i}-r)S_{i}\frac{\partial U}{\partial S_{i}}+\frac{1}{2}\sum_{i=1}^{d}\sum_{j=1}^{d}\rho_{i,j}\sigma_{i}\sigma_{j}S_{i}S_{j}\frac{\partial^{2}U}{\partial S_{i}^{2}}.$
@@
$+\sum_{i=1}^{d}v_{i,k}\frac{\partial U_{k}}{\partial\alpha_{i,k}}-v_{k}\sum_{i=1}^{d}(S_{i}f(v_{i,k}))\frac{\partial U_{k}}{\partial b_{k}}\}=0,$ (12)
with terminal condition
$U_{k}(S_{1},...,S_{d},b_{k},\alpha_{1,k},...,\alpha_{d,k},T)=(1-\gamma\psi)X_{k}(T)-\frac{\gamma}{2}X_{k}(T)^{2}.$
@@
Let $(S_{1},...,S_{d},b_{k},\alpha_{1,k},...,\alpha_{d,k},t)$ be denoted as $(S_{i},b_{k},\alpha_{i,k},t)$. Let $\Delta t$ be the timestep size such that $\Delta t\rightarrow 0$, then $\Delta S_{i}=S_{i}(t+\Delta t)-S_{i}(t),\Delta b_{k}=b_{k}(t+\Delta t)-b_{k}(t)$, and $\Delta\alpha_{i,k}=\alpha_{i,k}(t+\Delta t)-\alpha_{i,k}(t)$. Then by the optimality of the control and law of iterated expectation for $i=1,...,d$ and for each $k$ the value function of the auxiliary problem is given by:
$U_{k}=\sup_{v\in\mathcal{V}}\mathbb{E}[U_{k}(S_{i}+\Delta S_{i},b_{k}+\Delta b_{k},\alpha_{i,k}+\Delta\alpha_{i,k},t+\Delta t)].$ (13)
Let $\Delta U_{k}=U_{k}(S_{i}+\Delta S_{i},b_{k}+\Delta b_{k},\alpha_{i,k}+\Delta\alpha_{i,k},t+\Delta t)-U_{k}(S_{i},b_{k},\alpha_{i,k},t)$, then (13) can be written as
@@
Applying Ito’s Lemma on $\Delta U_{k}$, we get:
$\mathbb{E}[\Delta U_{k}]=\mathbb{E}\left[\frac{\partial U}{\partial t}\Delta t+\sum_{i=1}^{d}\frac{\partial U}{\partial S_{i}}\Delta S_{i}+\frac{1}{2}\sum_{i=1}^{d}\frac{\partial^{2}U}{\partial S_{i}^{2}}(\Delta S_{i})^{2}+\frac{\partial U}{\partial b_{k}}\Delta b_{k}+\sum_{i=1}^{d}\frac{\partial U}{\partial\alpha_{i,k}}\Delta\alpha_{i,k}\right],$
@@
$\mathbb{E}[\Delta S_{i}]$ $=(\mu_{i}-r+\sum_{k=1}^{K}g_{i,k}(v_{i,k}))S_{i}\Delta t,$
$\mathbb{E}[\Delta b_{k}]$ $=(rb_{k}-\sum_{i=1}^{d}v_{i,k}S_{i}f_{i,k}(v_{i,k}))\Delta t,$
$\mathbb{E}[\Delta\alpha_{i,k}]$ $=v_{i,k}\Delta t.$
The expectations of $\Delta S_{i},\Delta b_{k}$ and $\Delta\alpha_{i,k}$ are straight forward. The expectation of $(\Delta S_{i})^{2}$ requires more thought. Applying Ito’s lemma on $(\Delta S_{i})^{2}$, the expected value is given by:
$\mathbb{E}[(\Delta S_{i})^{2}]=\sum_{j=1}^{d}\rho_{i,j}\sigma_{i}\sigma_{j}S_{i}S_{j}\Delta t.$
Let $\mathcal{V}:=[v_{\min},v_{\max}]$ be the space of bounded trade rates $v$, such that there exists an optimal control $v_{i,k}^{*}$. Expanding (14) and evaluating $\mathbb{E}[\Delta S_{i}],\mathbb{E}[(\Delta S_{i})^{2}],\mathbb{E}[\Delta b_{k}],\mathbb{E}[\Delta\alpha_{i,k}]$ gives us:
$\left[\frac{\partial U_{k}}{\partial t}+\mathcal{L}U_{k}+rb_{k}\frac{\partial U_{k}}{\partial b_{k}}\right.$ $+\sup_{v\in\mathcal{V}}\left\{\sum_{i=1}^{d}\sum_{k=1}^{K}g_{i,k}(v_{i,k})S_{i}\frac{\partial U_{k}}{\partial S_{i}}\right.$
$\left.\left.+\sum_{i=1}^{d}v_{i,k}\frac{\partial U_{k}}{\partial\alpha_{i,k}}-\sum_{i=1}^{d}(v_{i,k}S_{i}f_{i,k}(v_{i,k}))\frac{\partial U_{k}}{\partial b_{k}}\right\}\right]\Delta t=0.$
Diving out the $\Delta t$ and bringing the supremum inside the expectation gives us the extended HJB equation with terminal condition
$U_{k}(S_{1},...,S_{d},b_{k},\alpha_{1,k},...,\alpha_{d,k},T)=(1-\gamma\psi)X_{k}(T)-\frac{\gamma}{2}X_{k}(T)^{2}.$
Remark 2.
The $K$ HJB equations of Lemma 3.1 are coupled through the optimal control.
@@
For the special case of $\kappa_{\tau}=0,\kappa_{p}=0$, i.e. there is no price impact, the optimal trade execution reduces to the optimal asset allocation problem. The analytical solution to the optimal asset allocation problem for time consistent mean-variance problem is known and has been solved in the linear-quadratic (LQ) sense [Zhou and Li 2000; Li and Ng 2000].
@@ 3.2 BSDE Formulation
A stochastic approach has an advantage in that the HJB can be extended to include multiple assets and multiple agents. This is computationally infeasible with the PDE approach even with the dimensional reduction techniques [Forsyth 2011]. This is because each PDE for variance and wealth requires a three dimensional grid even with one asset. Though dimensional reduction can be used, this is does not apply for the general performance conscious agent due to the averaging term. We drop the explicit dependence on $t$ to keep the notation concise. To solve (12), we formulate the BSDE for $U_{k}$ using the following lemma.
@@
If there exists the optimal control $v_{i,k}^{*}$, the BSDE that satisfies viscosity solution for (12) is given by:
@@
$\frac{\partial U_{k}}{\partial t}+\mathcal{L}_{k}U_{k}$ $=-rb_{k}\frac{\partial U_{k}}{\partial b_{k}}+\sum\limits_{i=1}^{d}v_{i,k}^{*}f_{i,k}(v_{i,k}^{*})\frac{\partial U_{k}}{\partial b_{k}}$
$-\sum\limits_{i=1}^{d}(v_{i,k}^{*})\frac{\partial U_{k}}{\partial\alpha_{i,k}}-\sum\limits_{i=1}^{d}\sum_{k=1}^{K}(g_{i,k}(v_{i,k}^{*}))S_{i}\frac{\partial U_{k}}{\partial S_{i}}.$
@@
The proof is from the direct application of the BSDE formulation given in [Pham 2014]. Let $x_{i,k}=\{S_{i},b_{k},\alpha_{i,k}\}$, for each agent $k$. Given the optimal control $v_{i,k}^{*}$ and a semi-linear PDE given by:
$\frac{\partial U_{k}}{\partial t}+\mathcal{L}U_{k}+\sum_{i=1}^{d}F(x_{i,k},v_{i,k}^{*},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)=0,$
@@
$dU_{k}=-\sum_{i=1}^{d}F(x_{i,k},v_{i,k}^{*},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)dt+\sum_{i=1}^{d}\sigma_{i}\frac{\partial U_{i,k}}{\partial x_{i,k}}dW_{i},$
from the semi-linear PDE we can express $-F(x_{i,k},v_{i,k}^{*},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)$ as:
$-\sum_{i=1}^{d}F(x_{i,k},v_{i,k}^{*},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)=\frac{\partial U_{k}}{\partial t}+\mathcal{L}U_{k}.$
@@
$-\sum_{i=1}^{d}F(x_{i,k},v_{i,k}^{*},U_{k},\frac{\partial U_{k}}{\partial x_{i,k}},t)$ $=-rb_{k}\frac{\partial U_{k}}{\partial b_{k}}+\sum\limits_{i=1}^{d}v_{i,k}^{*}f_{i,k}(v_{i,k}^{*})\frac{\partial U_{k}}{\partial b_{k}}$
$-\sum\limits_{i=1}^{d}(v_{i,k}^{*})\frac{\partial U_{k}}{\partial\alpha_{i,k}}-\sum\limits_{i=1}^{d}(g_{i,k}(\sum_{k=1}^{K}v_{i,k}^{*}))S_{i}\frac{\partial U_{k}}{\partial S_{i}}.$
The significance of the system of BSDEs is two fold. One, the system of BSDEs does not require us to compute the Hessian of the extended HJB equation. This a big advantage as it allows us to avoid computing the second derivative on $U$ since computation of the second derivative can be very costly for high dimensions. This reduces the computational complexity of the problem. The second significance of the system of BSDEs is that it allows us to solve (15) in a least squares sense which is a recent technique used in high-dimensional option pricing [Han and E 2016; E et al. 2017; Hure et al. 2020; Chen and Wan 2021; Na and Wan 2023]. This allows us to avoid using numerical solutions to the PDE and overcome the issue with dimensionality. Note, we can replace $X$ in (9) and (10) with $\hat{X}$ for the case where the agent is performance conscious, i.e. $\phi>0$.
@@ 3.3 Discretization of BSDE
$(S_{i})_{m}^{n+1}$ $=(1+(\mu_{i}-r+\sum_{k=1}^{K}g_{i,k}((v_{i,k})_{m}^{n}))\Delta t)(S_{i})_{m}^{n}+\sigma_{i}(S_{i})_{m}^{n}(\Delta W_{i})_{m}^{n},$ (16)
@@
The number of shares in stocks, $\alpha(t)$, is updated using (LABEL:eq:trade_flow_agent). Thus the discretized dynamics of $\alpha(t)$ is given by:
@@
Let $\hat{\mathcal{V}}$ be the discretization of $\mathcal{V}$ into $C\in\mathbb{Z}^{+}$ discrete points. Equations (16), (17), and (18) are simulated forward in time over a subset of discretized control values $\hat{\mathcal{V}}\subset\mathcal{V}$. The discrete set of control points does not need to be too dense and we restrict it to the discrete set shown in [Forsyth 2011].